Ho: μ1 - μ2 = 0

Ha: μ1 - μ2 ≠ 0

When testing above hypotheses the test statistic t was found to be 5.15. If the degrees of freedom = 40, then, the p-value for this test would be

Question options:

a) less than zero

b) between 0 and 0.5

c) less than .005

d) greater than 1
Answer #1

One sample t test value can be calculate as follows

t = 5.15

n = 40

Looking for the value t = 5.15 across n-1 = 40 - 1 = 39 degree of freedom in t table.

For two tailed test we get the value as P value = < 0.005

Part c) is correct.